In a pressure cooker, the potato is all cooked when approximately 49% of the cooking water has been evaporated. How much energy would be required for this, if initially there was 2.5L of cooking water at 22C ? Neglect the mass of potato in your calculation and provide answer in unit of MJ (mega joule), round it off to 2 significant figures.
Answer #1

Hi this is a two step process

1 L = 1000 mL = 1000 g (since water density = 1g/mL) ==>2.5 L = 2500 g;

water specific heat capacity = 4.18 J/g.°C;

1.

To heat the water from 22°C to boiling at 100°C

q1 = mcΔT = 2500g x 4.18 J/g.°C x (100 -22) °C = 815100 J

2.

To vaporize the water to steam at 100°C

49% of 2.5 L = 49/100 x 2.5 = 1.225 L = 1225 g is vaporized; enthalpy of vaporization = 2250 J/g

q2 = mcΔT = 1225 g x 2250 J/g = 2756250 J

Total heat energy required q = q1 + q2 = 815100 J + 2756250 J = 3571350 J ( 1 MJ = 100000 J)

3571350 J / 106 = 3.57135 MJ ~ 3.6 MJ

Answer: 3.6 MJ

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")